# Topology-Driven Quantum Architecture Search Framework

### v1

@@ 2.2.3 Quantum Neural Network for quantum state classification
Quantum Machine Learning (QML) is an advanced field that integrates quantum computing with machine learning, which aims to leverage the unique properties of quantum mechanics to enhance data analysis efficiency. In this study, we employ a Quantum Neural Network (QNN) to address a binary quantum state classification problem. We utilize the quantum entanglement dataset provided in Ref. [46], which consists of quantum states characterized by different levels of Concurrence Entropy (CE). This dataset is utilized for training the QNN, which comprises an embedding layer, PQC, and a measurement layer. The embedding layer is implemented by using angle encoding with RX rotation gates to map classical information to quantum states. The PQC component, searched by the QAS algorithm, processes the quantum states that contain the information from the dataset. Finally, the measurement layer measures the quantum states through Pauli-Z measurements to obtain classical bits, which are then processed by a multilayer perceptron (MLP) to provide the binary classification result.
@@ 3.1 TD-QAS Framework
Figure 1: Comparison of Standard QAS and Topology-Driven QAS Framework.
Standard QAS algorithms explore a vast search space that encompasses both circuit topology information (e.g., qubit involvement) and gate-type information. For instance, a fundamental element of a quantum circuit, such as $[R_{x}-0]$, specifies that the $R_{x}$ gate acts on qubit 0. Consequently, an entire circuit can be represented as a sequence of these elements, such as [$R_{x}-0$, $R_{z}-2$, $YY-0$, $XX-1$, $R_{y}-1$, $ZZ-0$, $R_{z}-2$], as illustrated in Figure 1 (a). However, this approach significantly amplifies the search difficulty of QAS. Specifically, heuristic search strategies and evaluation mechanisms must concurrently explore and optimize both topology and gate-type information within an extremely large search space, leading to substantial computational complexity and inefficiency. NAS has demonstrated that high-performance CNNs remain largely unaffected by the random replacement of specific operations [45]. If a similar phenomenon could be observed in QAS, separating the search for topology and gate-types would still enable the discovery of high-performance quantum circuits. This method can significantly reduce the search space by eliminating the need to search for gate configurations within low-performance topologies.
Through extensive numerical simulations, as discussed in Section 4.1, we found that high-performance quantum circuits exhibit only slight performance changes when quantum gate types are randomly replaced, provided that the circuit topology remains unchanged. This observation suggests that the key to searching for a high-performance VQA lies in determining an optimal circuit topology, followed by the selection of gate types. Therefore, we introduce the Topology-Driven Quantum Architecture Search (TD-QAS) framework, as illustrated in Figure 1 (b), which significantly reduces the search space size. Our framework first employs a topology-driven (TD) phase, which focuses on identifying high-performance topologies while keeping gate types fixed. Subsequently, the framework utilizes a gate tuning (GT) phase to adjust more suitable gate types to enhance circuit performance. To circumvent the substantial computational complexity associated with training the QAS of tuning phase from scratch, the tuning phase inherits parameters from the TD-QAS phase, enabling rapid convergence with only a few additional training epochs.
3.2 Implementation of TD-QAS Framework
To realize the TD-QAS framework, we define a topology representation that constructs a dedicated topology search space and establish an evaluation strategy tailored specifically for topology. In our framework, topology is defined by two types of placeholders: single placeholders and double placeholders. These placeholders represent the connectivity and arrangement of gate within the quantum circuit. As illustrated in Figure 1(b), a topology can be represented as [$single-0$, $single-2$, $double-0$, $double-1$, $single-1$, $double-0$, $single-2$]. The simplest method to represent topology employs a single placeholder, allowing any type of quantum gate at a given position. However, this approach is insufficient for accurately expressing topology. Moreover, we do not design a three-qubit placeholder or higher, as most QAS methods predominantly focus on single and double -qubit gates. Based on these considerations, defining two types of placeholders is both practical and effective.
To facilitate the search for topology, it is essential to establish an evaluation method specific to topology. The most straightforward approach to evaluate a topology’s performance would involve exhaustively enumerating or randomly generating a large number of quantum circuits and using their performance metrics as proxies for the topology’s effectiveness. However, this method is computationally prohibitive, especially as the size of the search space grows. We introduce a resource-efficient evaluation strategy termed the Topology Instantiation Evaluation Method. Specifically, we create a topology instance by replacing single and double placeholders with predesignated single-qubit gates (such as $R_{X}$) and double-qubit gates (such as $XX$ or $CNOT$). The performance of this topology instance serves as an approximation of the topology’s performance. Given that gate types have a relatively limited impact on the performance of high-performance circuits within QAS, this method provides a reasonable and efficient solution to evaluate topology without incurring substantial computational complexity. In our numerical simulations, the topology is directly represented as a specific topology instance. Section 4.2 will verify the accuracy of the topology instantiation evaluation method in reflecting the true performance of the topology.This validation ensures that our evaluation strategy reliably estimates the effectiveness of different topologies, thereby supporting the efficiency gains of the TD-QAS framework.
3.3 Reduction of Search Space
In this subsection, we analyze how the TD-QAS framework effectively reduces the search space, providing an intuitive understanding of its impact. First, the topology search space is primarily determined by the number of placeholder types. The gate-type search space is determined by the searched topology and the number of quantum gates available, which typically defined by manual design or hardware constraints. Given that the number of selectable quantum gates is usually limited, it is feasible to exhaustively enumerate all possible quantum circuits within the gate-type search space for small-scale problems.
To illustrate the efficacy of the TD-QAS framework in reducing the search space, consider a three-qubit quantum circuit consisting of seven quantum gates ($N_{\text{qubit}}=3$, $N_{\text{layer}}=7$) with ring connectivity among all qubits. Due to the ring connectivity, there are $N_{\text{qubit}}$ possible positions for placing either single-qubit and two-qubit gates. Suppose we define $N_{\text{single}}=3$ for the single-qubit gates set $\{R_{x},R_{y},R_{z}\}$ and $N_{\text{double}}=3$ for the two-qubit gates set $\{XX,YY,ZZ\}$. In classical QAS, the number of possible operations per layer is 18 $(N_{\text{qubit}}\cdot(N_{\text{single}}+N_{\text{double}}))$, and the search space size is $18^{7}$ $(\left[N_{\text{qubit}}\cdot(N_{\text{single}}+N_{\text{double}})\right]^{N_{\text{layer}}})$. the TD-QAS framework employs two types of placeholders, placeholders of each layer have 6 $(2\cdot N_{\text{qubit}}$) possible combination. Thus, the topology search space size becomes $6^{7}$ $(|(N_{\text{qubit}}\cdot 2)^{N_{\text{layer}}}|)$. The gate-type search space is $3^{7}$ $(|N_{\text{single}}^{x}\cdot N_{\text{double}}^{7-x}|)$, where $x$ represents the number of single placeholder in the topology. The TD-QAS framework decouples the operation search space into a topology search space and a gate-type search space, reducing it from $18^{7}$ to $6^{7}$ and $3^{7}$, the corresponding search space reduction ratio is nearly $2\times 10^{3}$, which calculation formula is (6). In this paper, we use reduction ratio to evaluate the compression efficiency of our framework.
$Reduction\ Ratio=\frac{S_{\text{proposed}}}{S_{\text{traditional}}}$ (6)
3.4 Quantum Computational Costs
In this work, we introduce the TD-QAS framework, which is characterized by exceptionally low computational complexity. For QAS, the foremost consideration is quantum computational complexity. Consequently, we utilize quantum computational costs to evaluate the computational overhead of QAS algorithms [35]. Quantum computational costs quantify the estimated time required to execute all quantum circuits generated by QAS algorithms on actual devices. The execution time $T$ for an individual quantum circuit can be expressed as:
$T=T_{p}+(circuit\ depth)\cdot T_{G}+T_{M}$ (7)
$T_{P}$ is the time to prepare the initial state, $T_{g}$ is the average duration of a quantum gate, and $T_{M}$ is the time needed to measure the qubits. where $T_{p}+T_{M}$ is 1 microsecond, and the average gate time $T_{G}$ is set to 0.01 microseconds. Circuit depth refers to the maximum length of the directed path from input to output in the quantum circuit, which is directly related to the coherence time required to complete the algorithm. In our numerical simulations, we will compare the quantum computational costs between TD-QAS framework and original QAS algorithms, aiming to evaluate their respective efficiencies.
4 Numerical Simulation
In this section, we conduct a comprehensive set of numerical simulations to evaluate the proposed TD-QAS framework across three representative tasks: estimation of ground state energies for molecular systems, solving the MaxCut problem on 100 Erdős-Rényi (ER) graphs, and performing an 8-qubit quantum state classification task.
We first validated that high-performance quantum circuits are only slightly affected in performance when quantum gate types are randomly replaced while maintaining the circuit topology. This indicates that prioritizing the search for circuit topology before fine-tuning gate types is a reasonable approach. Secondly, we validated that the topology instantiation evaluation method can accurately assess circuit topology performance, offering an evaluation strategy that enables QAS algorithms to efficiently search for circuit topologies with low complexity. Lastly, we verify the superiority of our TD-QAS framework by integrating it with two representative QAS algorithms: QuantumSupernet [23] and DQAS [24]. Our results show that the TD-QAS framework enhances QAS algorithms by identifying higher-performance quantum circuits with lower computational complexity compared to their standard counterparts. Furthermore, we verified that the TD-QAS framework can be easily deployed within existing QAS algorithms, highlighting its versatility and potential to advance the scalability of QAS methodologies.
The purpose of this section is to validate the core hypothesis of the TD-QAS framework, which posits that quantum circuit topology plays a dominant role over the type of quantum gates in determining VQA performance. This hypothesis is based on our preliminary observations, which suggest that for high-performing circuits, keeping the topology fixed while randomly modifying the quantum gate-type does not significantly impact overall performance. To verify the generality of this phenomenon, we conducted a series of numerical simulations across three different tasks.
We first randomly generated 1,000 original quantum circuits for three distinct tasks. Each quantum circuit consisted of either 30, 35, or 50 quantum gates chosen from the native gate set $\{H,R_{x},R_{y},R_{z},XX,YY,ZZ\}$, with qubit connectivity arranged in a ring connectivity. For each original circuit, we created 10 variants by modifying the gate-type. Specifically, we selected one-fifth of the gates in the circuit and replaced them with other gates of the same type (i.e., single-qubit gates replaced by other single-qubit gates), which ensured that the topology remained unchanged. An original-variant set was constructed to verify whether the variants of a high-performance original circuit exhibit reduced performance. After generating these original-variant pairs, each circuit was initialized with parameters five times and trained until convergence. We then selected the best performance from the five training runs as the true performance. This procedure produced data pairs consisting of the original circuit performance and the corresponding performance of its variants [$p_{i}$, $v_{i}^{j}$], where $i$ represents the original circuit ($i\in[1,1000]$) and $j$ represents the variant circuit ($j\in[1,10]$). To quantify the performance differences between each original circuit and its variants, we computed the mean squared error (MSE) between them, yielding pairs of original performance and performance difference [$p_{i}$, $d_{i}$] as defined:
$d_{i}=\frac{1}{n}\sum_{i=1}^{n}\left(p_{i}-p_{i}^{j}\right)^{2}\quad i\in[1,500],j\in[1,10]$ (8)
These data were plotted as blue dots in Figure 2. To further illustrate the differences at varying performance levels, the range of $p_{i}$ was divided into multiple intervals (100, 100, and 50 intervals respectively), and the average $d_{i}$ for each interval was calculated and represented by the green line.
Figure 2: Simulation results for the three tasks: (a) VQE for ground state energy estimation, (b) VQA for MaxCut problem on Erdős-Rényi(ER) graphs, and (c) QNN for 8-qubit quantum state classification. The horizontal axis represents the original circuit’s performance $p_{i}$, and the vertical axis represents the performance difference $d_{i}$. Blue dots show individual performance differences between original circuits and their variants, while the green line indicates the mean difference across different performance levels.
The results from Figure 2 show a consistent trend: as the performance of the original circuit improves, the influence of gate modifications on overall performance diminishes. This trend was consistently observed across all three tasks, highlighting the generality and robustness of our findings. These results strongly support the core hypothesis of the TD-QAS framework.
4.2 Topology Instantiation Evaluation Method
The TD-QAS framework proposes that QAS should first search for topology. However, existing QAS methods lack a dedicated method for evaluating topology. The most straightforward way to evaluate a topology’s performance is to exhaustively assess all potential quantum circuits or a large number of circuits and then use their average performance as a proxy for the topology’s performance. Unfortunately, this approach is impractical due to its high complexity. As observed in Section 4.1, we find that random gate substitutions have minimal impact on performance for high-performance circuits. Therefore, we propose a highly resource-efficient evaluation strategy called the topology instantiation evaluation method. This method represents the topology using topology instances, which replace the corresponding placeholders with specific quantum gates. To verify the effectiveness of this method, we calculated the correlation between the approximate true performance and the instance performance calculated by our method.
We first generated 100 circuit topologies. For each topology, we generated 100 quantum circuits by randomly selecting quantum gates from the set $\{H,R_{x},R_{y},R_{z},XX,YY,ZZ\}$, arranged in a ring connectivity. We then calculated the performance obtained from two evaluation methods. For each topology, we initialized the parameters for the corresponding 100 circuits and trained them to obtain their average performance, which served as the approximate true performance, denoted as $y$. In the our proposed method, we replaced the single-qubit and two-qubit placeholders with $R_{x}$ gates and $XX$ gates, respectively, generating one instance for each topology. Similarly, we calculated the performance of these instances, denoted as $y^{\prime}$. Finally, we calculated the Pearson correlation coefficient between the true performance, $y$, and the instance performance, $y^{\prime}$, as defined in (9):
$r=\frac{\sum{(y_{i}-\bar{y})(y^{\prime}_{i}-\bar{y^{\prime}})}}{\sqrt{\sum{(y_{i}-\bar{y})^{2}}\cdot\sum{(y^{\prime}_{i}-\bar{y^{\prime}})^{2}}}}$ (9)
The numerical simulation results indicated correlation values of 0.68, 0.59, and 0.62 for the three tasks. We present the results for one task in Figure 3. For comparison, the widely used performance predictor achieved a correlation of 0.71 [25], while it required extensive complexity to prepare training datasets. Thus, our method can provide a reasonably accurate assessment of topology’s performance with low complexity.
Figure 3: Correlation of true performance and instance performance on the state classification task.
4.3 TD-QAS framework
@@
In the previous sections, we discussed the challenges faced by existing QAS algorithms, including the vast search spaces that often lead to local optima and excessive quantum computational costs. To tackle these challenges, we propose the TD-QAS framework, which aims to enhance the efficiency of QAS by decomposing the operational search space into separate topology and gate-type search spaces. To thoroughly assess the effectiveness of the TD-QAS framework, we compare several metrics from both the TD and GT phases with those of standard QAS algorithms. These metrics include the size of the search space, performance in specific tasks, circuit properties, and overall quantum computational costs. In the following subsections, we apply the TD-QAS framework to two representative QAS algorithms and analyze their performance.

### v2

@@ 2.2.3 Quantum Neural Network for Quantum State Classification
Quantum Machine Learning (QML) is an advanced field that integrates quantum computing with machine learning, aiming to exploit the unique properties of quantum mechanics to enhance data analysis efficiency. In this study, we employed a Quantum Neural Network (QNN) to address a binary quantum state classification problem. We utilized the quantum entanglement dataset provided in Ref. [46], which consists of quantum states characterized by different levels of Concurrence Entropy (CE). This dataset was utilized for training the QNN, which comprises an embedding layer, PQC, and a measurement layer. The embedding layer was implemented using angle encoding with $R_{x}$ rotation gates to map classical information to quantum states. The PQC component, searched by the QAS algorithm, processes the quantum states that contain the information from the dataset. Finally, the measurement layer extracts classical information from quantum states using Pauli-Z measurements, which are processed by a multilayer perceptron (MLP) to provide the binary classification result.
@@ 3 Method
3.1 Motivation and overview
Figure 1: Comparison between standard QAS and Topology-Driven QAS (TD-QAS). (1) In standard QAS, the search must jointly explore both topology and gate types, causing a large and complex search space. (2) TD-QAS decouples the search into two sequential stages, which first identifies optimal topologies using symbolic placeholders, and then assigns gate types during a separate tuning phase. Here, (a) and (c) represent complete quantum circuits with both topology and gate-type information, while (b) depicts a topology-only representation using placeholders.
Standard QAS algorithms explore a vast search space that encompasses both circuit topology information, such as qubit involvement, and gate-type information. For instance, a fundamental element of a quantum circuit, such as $[R_{x}0]$, specifies that the $R_{x}$ gate acts on qubit 0. Consequently, an entire circuit can be represented as a sequence of these elements, such as [$R_{x}0$, $R_{z}2$, $YY0$, $XX1$, $R_{y}1$, $ZZ0$, $R_{z}2$], as illustrated in Figure 1 (a). However, this approach significantly increases the search difficulty in QAS. Specifically, heuristic search strategies and evaluation mechanisms must simultaneously explore both topology and gate-type information. This joint exploration across an extremely large search space results in substantial computational complexity and inefficiency. NAS has demonstrated that high-performance CNNs remain unaffected by the random replacement of specific operations [47]. If a similar phenomenon occurs in VQA, decoupling topology and gate-type searches could enhance QAS in discovering high-performance quantum circuits. This method can significantly reduce the search space by eliminating the need to search for gate configurations within low-performance topologies.
Extensive numerical simulations, as discussed in Section 4.1, reveal that high-performance quantum circuits exhibit only slight performance changes when quantum gate types are randomly substituted, provided that the circuit topology remains unchanged. Remarkably, 76% of the quantum circuits exhibited no significant change in performance after their gate types were altered in noise scenario. This observation indicates that identifying an optimal circuit topology is the primary step in searching for a high-performance VQA, followed by gate-type selection. Therefore, we introduce the Topology-Driven Quantum Architecture Search (TD-QAS) framework, as illustrated in Figure 1 (b), which significantly reduces the search space size. Our framework first employs a topology-driven (TD) phase, which prioritizes the identification of high-performance topologies. Subsequently, the framework utilizes a gate tuning (GT) phase to adjust the suitable gate types to enhance circuit performance. To circumvent the substantial computational complexity associated with training the QAS of the tuning phase from scratch, the tuning phase inherits parameters from the TD-QAS phase, enabling rapid convergence with only a few additional training epochs.
3.2 Implementation of TD-QAS framework
Presentation and evaluation strategy collectively enable topology search within QAS algorithms. We emphasized topology representation and a dedicated search space. At the algorithmic description level, circuit topology is defined as a sequence of placeholders, each representing a potential position for a quantum gate. The type of placeholder depends on the number of qubits the gate acts on. As shown in Figure 1(b), a topology might be represented as [$single$0, $single$2, $double$0, $double$1, $single$1, $double$0, $single$2], where single placeholders correspond to single-qubit gates, and double placeholders correspond to two-qubit gates. This configuration explicitly defines the topology search space for a task with a fixed number of qubits and quantum gates.
To facilitate the search for topology, it is essential to establish an evaluation method specific to topology. The straightforward approach to evaluate a topology’s performance is to exhaustively enumerate or randomly generate numerous quantum circuits, using their performance metrics as proxies for the topology’s effectiveness. However, this method becomes computationally prohibitive, especially as the search space expands. We introduce a resource-efficient topology evaluation strategy called the Topology Instantiation Evaluation Method. In this approach, a topology instantiation is created by replacing single- and double-qubit placeholders with predefined gates, such as $R_{X}$ for single-qubit and $XX$ (or $CNOT$) for two-qubit operations. The performance of a topology instantiation provides an approximation of the topology’s true performance. Considering the relatively limited impact of gate types on circuit performance within QAS, this method provides a reasonable and efficient solution to evaluate topology without incurring substantial computational complexity. Section 4.2 verifies the accuracy of the topology instantiation evaluation method in reflecting the true performance of the topology. This validation ensures that our evaluation strategy reliably estimates the effectiveness of different topologies, thereby supporting the efficiency of the TD-QAS framework.
The proposed topology representation and evaluation strategy collectively enable topology search within QAS algorithms. This topology search readily integrates into mainstream QAS algorithms without requiring any modifications to their core search mechanisms. Specifically, when the native gate set is limited to a single-qubit gate and one type of two-qubit gate, mainstream QAS algorithms transition from searching for quantum circuits to focusing on topology. This design aligns well with our proposed Topology Instantiation Evaluation method. Consequently, at the level of practical implementation, all topology representations can be evaluated through their instantiations. After completing topology search, the TD-QAS framework proceeds to the gate-type search phase. The size of the gate-type search space depends on both the selected topology and the native gate set. The core mechanisms, including the evaluation strategy and search procedure, remain unchanged and are omitted here for brevity. In summary, the TD-QAS framework can be implemented on existing QAS methods with minimal modifications.
3.3 Reduction of search space
In this subsection, we analyze how the TD-QAS framework effectively reduces the search space, offering an intuitive perspective of its impact. First, the topology search space is typically determined by the number of placeholder types and the properties of the quantum circuit, such as the number of qubits and gates. The gate-type search space is determined by the searched topology and the native gate set, which is typically defined by manual design or hardware constraints. To illustrate the efficacy of the TD-QAS framework in reducing the search space, consider a three-qubit quantum circuit consisting of seven quantum gates ($N_{\text{qubit}}=3$, $N_{\text{gate}}=7$) with ring connectivity among all qubits. We assume the qubit connectivity is ring connectivity, and there are $N_{\text{qubit}}$ possible positions for placing either single-qubit or two-qubit gates. Suppose we set the native gate set $A=\{R_{x},R_{y},R_{z},XX,YY,ZZ\}$, thus $|A[\text{single}]|=3$ and $|A[\text{double}]|=3$, where $|A[\text{single}]|$ denotes the number of single-qubit gates in the native gate set. In classical QAS, the number of possible operations per layer is 18 $(N_{\text{qubit}}\cdot(|A[\text{single}]|+|A[\text{double}]|))$, and the search space size is $18^{7}$ $\left[N_{\text{qubit}}\cdot(|A[\text{single}]|+|A[\text{double}]|)\right]^{N_{\text{gate}}}$. The TD-QAS framework employs two types of placeholders, with placeholders in each layer having 6 $(2\cdot N_{\text{qubit}})$ possible combinations. Thus, the topology search space size becomes $6^{7}$ $(|(N_{\text{qubit}}\cdot 2)^{N_{\text{gate}}}|)$. The gate-type search space is $3^{7}$ $(|A[\text{single}]|^{x}\cdot|A[\text{double}]|^{7-x})$, where $x$ represents the number of single placeholders in the topology. The TD-QAS framework decouples the operation search space into a topology search space and a gate-type search space, reducing it from $18^{7}$ to $6^{7}$ and $3^{7}$. The corresponding search space reduction ratio is approximately $2\times 10^{3}$, and the calculation formula is provided in (6). In this study, we use the reduction ratio $R$ to evaluate the compression efficiency of our framework.
$R=\frac{S_{\text{proposed}}}{S_{\text{traditional}}}.$ (6)
3.4 Quantum computational costs
In this work, we introduce the TD-QAS framework, which is characterized by exceptionally low computational complexity. In QAS, quantum computational complexity is the primary consideration. Consequently, we utilize quantum computational costs to evaluate the computational overhead of QAS algorithms [36]. Quantum computational costs estimate the execution time for all quantum circuits generated by QAS algorithms on actual devices. The execution time $T$ for an individual quantum circuit can be expressed as:
$T=T_{p}+(circuit\ depth)\cdot T_{G}+T_{M},$ (7)
where $T_{P}$ is the time to prepare the initial state, $T_{g}$ is the average duration of a quantum gate, and $T_{M}$ is the time required to measure the qubits. Here, $T_{p}+T_{M}$ is 1 microsecond, and the average gate time $T_{G}$ is set to 0.01 microseconds. Circuit depth represents the maximum length of the directed path from input to output in the quantum circuit, directly influencing the coherence time required to complete the algorithm. In our numerical simulations, we will compare the quantum computational costs between TD-QAS framework and original QAS algorithms, aiming to evaluate their respective efficiencies.
4 Numerical simulation
In this section, we conduct extensive numerical simulations to evaluate the proposed TD-QAS framework across three representative tasks both noiseless and noisy scenarios. Specifically, three representative tasks include estimating ground state energies for molecular systems, solving the MaxCut problem on 100 Erdős-Rényi (ER) graphs with 10 nodes, and performing an 8-qubit quantum state classification task. In the noisy scenario, we adopted depolarizing and bit-flip noise models, with parameters set to 0.01 for single-qubit depolarizing, 0.001 for two-qubit depolarizing, and 0.01 for bit-flip noise.
In the following sections, we first validated that high-performance quantum circuits experience only minor performance variations when quantum gate types are randomly replaced, provided the circuit topology remains unchanged. Thus, prioritizing the search for circuit topology before fine-tuning gate types is a reasonable approach. Secondly, we validated that the topology instantiation evaluation method effectively assess circuit topology performance, offering an evaluation strategy that enables QAS algorithms to efficiently search for low-complexity circuit topologies. Finally, we verify the superiority of our TD-QAS framework by integrating it with two representative QAS algorithms, which is QuantumSupernet [23] and DQAS [24].
This subsection aims to validate the core hypothesis of the TD-QAS framework, which posits that quantum circuit topology has a greater influence on VQA performance than quantum gate types. This hypothesis is motivated by our preliminary observations, suggesting that for high-performing circuits, maintaining a fixed topology while randomly modifying the quantum gate-type has minimal impact on overall performance. To verify the generality of this phenomenon, we conducted a series of numerical simulations. Specifically, we first describe the design and settings of numerical simulations. Thereafter, we validate the core hypothesis using a VQE task under both noiseless and noisy scenarios. Finally, we extend the evaluation to additional tasks, including the MaxCut problem and quantum state classification, to further demonstrate the robustness of the hypothesis across different task.
Figure 2: Verification of the core hypothesis on the VQE task under noiseless (left) and noisy (right) conditions. The horizontal axis shows the original circuit’s performance, while the vertical axis indicates the performance difference between the original circuits and their gate-modified variants. Blue dots represent individual performance differences; the green line denotes the average difference across performance levels.
We outline the numerical simulation process, comparing the performance of the original quantum circuits with their gate-modified variants. Specifically, we first randomly generated 1,000 original quantum circuits. Each quantum circuit consisted of multiple quantum gates chosen from the native gate set $\{R_{x},R_{y},R_{z},XX,YY,ZZ\}$, with qubit connectivity arranged in a ring connectivity. For each original circuit, we generated 10 variants by randomly modifying 20% of the gate types, ensuring that each replaced gate remained within the same class (i.e., single-qubit gates replaced by other single-qubit gates), thereby preserving the topology. We selected the best performance from the five training runs as the true performance. This performance dataset consist of pairs, each containing the original circuit performance and that of its corresponding variants [$p_{i}$, $v_{i}^{j}$], where $i$ represents the original circuit ($i\in[1,1000]$) and $j$ represents the variant circuit ($j\in[1,10]$). To quantify the performance differences between each original circuit and its variants, we computed the mean squared error (MSE), yielding pairs of original performance and performance difference [$p_{i}$, $d_{i}$] as defined (8):
$d_{i}=\frac{1}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}\left(p_{i}-p_{i}^{j}\right)^{2},$ (8)
We first evaluated this hypothesis on a VQE task (6-qubit TFIM) under both noiseless and noisy scenarios. Specifically, both the original circuits and their variants were configured to contain 35 quantum gates. The numerical simulation results are visualized in Figure 2, where each blue dot represents a circuit-variant pair, and the green curve shows the average $d_{i}$ over performance intervals (100 and 10 bins, respectively). To better observe this trend under the noisy scenario, we excluded extreme outliers with performance below -5. In both noiseless and noisy scenarios, we observe a consistent trend: as circuit performance improves, the impact of gate-type changes diminishes. Remarkably, in the noisy VQE task, 76% of circuits exhibited minimal performance change (MSE $<$ 0.1) after gate types were altered. We attribute this phenomenon to the influence of noise, which overwhelms the relatively small effects caused by changes in gate types. These findings provide strong support for the core hypothesis.
To further assess the robustness of the hypothesis, we extended the simulation protocol to two additional tasks: solving the MaxCut problem and performing quantum state classification. For each task, the original circuits and their variants contained 35 or 50 quantum gates, and their performance was evaluated under noiseless scenarios. The results, shown in Figure 3, demonstrate a similar trend to that observed in the VQE experiments, reinforcing the generality of our core hypothesis across different tasks.
Figure 3: Verification of the core hypothesis on solving the MaxCut problem (left) and quantum state classification task (right). The horizontal axis represents the original circuit’s performance, and the vertical axis represents the performance difference between the original circuits and their gate-modified variants. Blue dots represent individual performance differences, while the green line shows the average difference across different performance levels.
4.2 Topology instantiation evaluation method
Existing QAS methods lack a dedicated approach for evaluating topology. The ideal method to evaluate a topology’s performance is to exhaustively analyze all potential quantum circuits or numerous circuits, and use their average performance as a proxy for topology effectiveness. However, this approach is impractical owing to its high computational complexity. As observed in Section 4.1, we found that random gate substitutions have minimal impact on performance for high-performance circuits. Therefore, we propose a highly resource-efficient topology evaluation method called the Topology Instantiation Evaluation. This method represents the topology using topology instantiation, where the corresponding placeholders are replaced with specific quantum gates. To validate the effectiveness of this method, we computed the correlation between the approximate true performance and the performance of the topology instantiation.
We first generated 100 circuit topologies. For each topology, we created 100 quantum circuits by randomly selecting quantum gates from the set $\{R_{x},R_{y},R_{z},XX,YY,ZZ\}$. For each topology, we initialized parameters for 100 corresponding circuits, trained them, and use their average performance as the approximate true performance, denoted as $\bar{y}$. Subsequently, for each of topologies, we replaced the single-qubit and two-qubit placeholders with $R_{x}$ gates and $XX$ gates, respectively. This process generated one instantiation circuit for each topology. Similarly, we calculated the performance of these instantiations, denoted as $y^{\prime}$. Finally, we computed the Pearson correlation coefficient between the true performance, $y$, and the instantiation performance, $y^{\prime}$.
The numerical simulations yielded correlation values of 0.68, 0.59, and 0.62 for the three tasks. We present the results for one task in Figure 4. For comparison, the widely used performance predictor achieved a correlation of 0.71 [25], although it required considerable complexity to prepare the training datasets. Thus, our method provides a reasonably accurate assessment of topology’s performance while maintaining significantly lower complexity.
Figure 4: Correlation between true performance and instantiation-based approximation in the state classification task. Each point represents a topology, where the x-axis indicates the true performance (averaged over 100 gate-randomized circuits), and the y-axis shows the performance estimated by the Topology Instantiation Evaluation Method.
To further evaluate the robustness of the proposed Topology Instantiation Evaluation method, we conducted an additional experiment on the VQE task. In particular, we evaluated two alternative combinations of single- and double-qubit gates ($R_{y}$ with $YY$ and $R_{y}$ with $XX$), measuring the instantiation performance and correlating it with the true average performance. The correlation values obtained with these gate sets were 0.68 and 0.65, respectively, which are closely aligned with the correlation of 0.68 observed when using the $R_{x}$ and $XX$ gates. These results indicate that the accuracy of the proposed evaluation method remains unaffected by the specific gate types used for instantiation.
@@ 4.3 TD-QAS framework
In the previous sections, we discussed the challenges faced by existing QAS algorithms, including the vast search spaces and excessive computational costs. Therefore, we propose the TD-QAS framework, which aims to enhance the efficiency of QAS by decomposing the operational search space into separate topology and gate-type search spaces. To assess the effectiveness of the TD-QAS framework, we compare several metrics from both the TD and GT phases with those of standard QAS algorithms. These metrics include the size of the search space, performance in specific tasks, circuit properties, and overall quantum computational costs. In the following subsections, we apply the TD-QAS framework to two representative QAS algorithms and analyze their performance.